OpenStudy (anonymous):

A pyramid has a square base of side x (in cm), and a height h (in cm) and its volume V is given by the formula V = 1/3x^2h. If the sum of the length of the base side and the height is 9, find the max volume. Give the dimensions of maximum volume.

OpenStudy (anonymous):

Use differentiation to find the first and second derivative.

OpenStudy (anonymous):

You have to get rid of h first though.

OpenStudy (anonymous):

If the sum of the length of the base side and the height is 9 <-----this already gives you an equation.

OpenStudy (anonymous):

Give an equation for that statement?

OpenStudy (anonymous):

so it would be \[x + h = 9\] and then you'd solve that for h, \[h = 9 - x\] then put that in for h in the original equation for V?

OpenStudy (anonymous):

And then get the first & second derivatives?

OpenStudy (anonymous):

Yep. replace h first before doing that.

OpenStudy (anonymous):

Then you can differentiate with respect to x.

OpenStudy (anonymous):

And you find the S.P's (stationary points) by letting the first derivative equal to zero.

OpenStudy (anonymous):

Find the x-values for the stationary points, then sub the x-value into the original equation. Once you do that, sub the x-values into the second derivative to see whether it's a maximum or minimum. If the second derivative is less than zero, it's a maximum if it's larger/greater than zero, it's a minimum.

OpenStudy (anonymous):

could you rewrite that equation for the volume. Could you draw it out so I can see it more accurately.

OpenStudy (anonymous):

Just to check my math so I don't mess it up, if \[h = 9 - x\] and I'm plugging that into \[V = 1/3x^2h\] it would be \[V = 3x^2-1/3x^3\] right?

OpenStudy (anonymous):

Is it \[\frac{1}{3}x^2h\] or \[\frac{1}{3x^2h}\]

OpenStudy (anonymous):

\[\frac{ 1 }{ 3 } x^2h\]

OpenStudy (anonymous):

Okay. \[\frac{ dV }{ dx }=\frac{ 18x-3x^2 }{ 3 }\] \[\frac{ d^2V }{ dx^2 }=-\frac{ 6x }{ 3 }\]

OpenStudy (anonymous):

@alaskamath

OpenStudy (anonymous):

At S.Ps dV/dx=0 \[18x-3x^2=0\] \[6x-x^2=0\] \[x(6-x)=0\] \[x=0, 6\] You can't have a length of zero so of course the length is 6. \[x=6\] Sub that into the original equation to find the volume.

OpenStudy (anonymous):

And if you sub in any positive x-value into the second derivative it's always negative. So it doesn't matter.

OpenStudy (anonymous):

So what was your equation for V after you subbed out h? Is it \[V = \frac{ 1 }{ 3 }x^2(9-x)\]?

OpenStudy (anonymous):

@Azteck

OpenStudy (anonymous):

Oh sorry about that.

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

sub x=6 into that.

OpenStudy (anonymous):

to find the volume.

OpenStudy (anonymous):

Okay, thank you!

OpenStudy (anonymous):

No worries.